### v4

# Explainable Artificial Intelligence for Autonomous Driving: A Comprehensive Overview and Field Guide for Future Research Directions

Abstract
Autonomous driving has achieved significant milestones in research and development over the last two decades. There is increasing interest in the field as the deployment of autonomous vehicles (AVs) promises safer and more ecologically friendly transportation systems. With the rapid progress in computationally powerful artificial intelligence (AI) techniques, AVs can sense their environment with high precision, make safe real-time decisions, and operate reliably without human intervention. However, intelligent decision-making in such vehicles is not generally understandable by humans in the current state of the art, and such deficiency hinders this technology from being socially acceptable. Hence, aside from making safe real-time decisions, AVs must also explain their AI-guided decision-making process in order to be regulatory compliant across many jurisdictions. Our study sheds comprehensive light on the development of explainable artificial intelligence (XAI) approaches for AVs. In particular, we make the following contributions. First, we provide a thorough overview of the state-of-the-art and emerging approaches for XAI-based autonomous driving. We then propose a conceptual framework that considers all the essential elements for explainable end-to-end autonomous driving. Finally, we present XAI-based prospective directions and emerging paradigms for future directions that hold promise for enhancing transparency, trustworthiness, and societal acceptance of AVs.
keywords:
Explainable artificial intelligence , autonomous driving , intelligent transportation systems , regulatory compliance , safety
[…]
1 Introduction
A survey of the American National Highway Traffic Safety Administration (NHTSA) reports that nearly 94% of road accidents are due to human errors [1]. These human-related mistakes are mainly classified as driver distraction, drunk or impaired driving, lack of attention, violation of the traffic rules, limited view of traffic conditions, and jay-walking pedestrians [2]. The lack of rule obedience, the increasing number of vehicles on roads, and improper road culture have, therefore, motivated officials, manufacturers, and legislators to make substantial improvements in transportation systems. There are growing research and development attempts to enhance safety and automation capability of AVs, prevent traffic accidents, and create a better road infrastructure. In this sense, the potential benefits of AVs are improved convenience, operational safety (especially for seniors and people with reduced mobility) [3], reduced $CO_{2}$ emissions [4], diminished transportation costs [5], improved reliability [6, 7], and reduced traffic density [8]. Intel’s report on the projected benefits of AVs estimates that deployment of this technology on roads will result in a reduction of 250 million hours of users’ commuting time per year and save more than half a million lives from 2035 to 2045, just in the USA [9]. While the potential impact
Figure 1: A canonical exemplar of explainable AI in autonomous driving: An autonomous vehicle provides a natural and intelligible explanation of its real-time decision to bystanders. Graphics adapted and modified from the source: [10].
and benefits of AVs in everyday life are promising, there is a major societal concern about functional safety of such vehicles. This issue, as a major drawback, originates mainly from reports of recent traffic accidents with the presence of AVs, primarily owing to their opaque and inappropriate decision-making [11, 12, 13, 14]. As AI approaches provide the foundation for real-time driving actions, there is an inherent need and expectation from consumers and regulators that AI-driven decisions of AVs should be explainable  [^1] We’ll use the terms explainable and interpretable interchangeably. (e.g., Figure 1) to build confidence in these vehicles. In a recent study, we have described the fundamental concepts and process steps associated with XAI-based autonomous driving [15]. In this paper, we extend the scope of that work by discussing the following research questions in depth:
  1. 1.
    Why is there a need for XAI in AVs technology?
  2. 2.
    What are the current trends and emerging AI technologies for explainable autonomous driving?
  3. 3.
    What are promising future XAI directions toward trustworthy, responsible, and publicly acceptable AVs?
With these questions in mind, our paper makes the following contributions:
  • We describe cross-disciplinary perspectives necessitating explainability in autonomous driving;
  • We provide a survey of state-of-the-art XAI-based investigations for autonomous driving;
  • We present a conceptual design framework for explainable end-to-end autonomous driving;
  • We propose future research directions on promising XAI approaches for autonomous driving.
The rest of the article is structured as follows. In Section 2, we provide background information and the factors triggering the need for the emergence of XAI in autonomous driving. We then describe the concept of explanations in autonomous driving by analyzing 1) cross-disciplinary perspectives necessitating explanations, 2) the role of various types of explanations for diverse explanation recipients, and 3) construction methodologies for explanations in Section 3. We present a brief overview of frontier AI approaches powering the real-time action decisions of AVs in Section 4. In Section 5, we provide a comprehensive survey of state-of-the-art investigations on explainable AI-based autonomous driving. Motivated by current limitations and trends delineated in these works, we present a general design framework for explainable autonomous driving and show key components of such a framework in Section 6. Finally, we describe potential challenges and a road map toward safety and explainability of next-generation AVs as future directions and sum up the article.
2 Background
2.1 AVs at a glance
AVs are intelligent vehicles capable of sensing their environment and mapping such sensing data to real-time driving decisions using an AI-based driving system. To discern, identify, and distinguish the objects in their operational surroundings, these vehicles fuse information from a variety of sensors that help make real-time driving decisions [16, 17]. The history of contemporary AVs goes back to 1988, when ALVINN (Autonomous Land Vehicle In a Neural Network), the first neural network-powered self-driving vehicle taking camera images with a laser range finder, was able to produce control commands for the road-following task [18]. Current AVs deployed on road networks have different levels of automation based on their in-vehicle technologies and intelligent capabilities. SAE International (previously known as the Society of Automotive Engineers) has defined six levels of autonomous driving [19]: Level 0 - No automation (a human driver is responsible for all critical driving tasks); Level 1 - Driving assistance (a vehicle has automated driving support such as acceleration/braking or steering, but the driver is responsible for all other possible driving operations); Level 2 - Partial automation (Advanced Driving Assistance Systems (ADAS) operations such as steering and acceleration/braking are available in this level); Level 3 - Conditional automation (a vehicle has more advanced features such as object/obstacle detection and can carry out the majority of driving operations); Level 4 - High automation (a vehicle can fulfill all possible driving operations in a geofenced area); and Level 5 - Full automation (a vehicle can perform all driving operations in any likely scenario, and no human intervention is required).
In general, there are two main approaches to building autonomous driving systems in terms of their AI-based learning architecture: modular and end-to-end pipelines [13, 20]. The modular pipeline consists of four primary and interconnected modules categorized as perception, localization, planning, and control (Figure 2, a). Perception is the sensing of an operational environment defined as a combination of two tasks: road surface extraction and on-road object detection [21]. Information for the purpose of perception can be obtained via an integrated sensor suite, such as LIDAR, RADAR, visible spectrum cameras, and ultrasonic sensors [17, 22]. The process of localization enables an autonomous vehicle to accurately determine its position in the sensor model of the world [23, 24]. The most effective way to get the position of an autonomous car is to use satellite navigation-based systems. Among such systems, the Global Navigation Satellite System (GNSS) and its most popular type, Global Positioning System (GPS), is a universal sensor to determine the global location of a car [25]. Planning is a complex operation that integrates components of route planning (i.e., selection of a route in the road work from the initial point to the final destination) and behavior planning (such as interactions with other vehicles, and people that may be met on a trajectory). Finally, control of an autonomous vehicle is the appropriate execution of chosen actions. In modern AVs, control is typically carried out through the ADAS software.
Overall, the modular pipeline leverages various sensor suites and algorithms for each module. While being comprised of standalone components makes the modular system more explainable and debuggable, such an architecture propagates errors to the next component, and thus, the overall pipeline error becomes cumulative [20, 26, 27].
In contrast to a modular pipeline, end-to-end autonomous driving has recently emerged as a paradigm
Figure 2: Modular vs. end-to-end autonomous driving. In the modular pipeline, the described four operations are carried out subsequently to produce control commands while end-to-end driving directly inputs raw sensor data and produces control commands.
shift in the design and development of AVs. End-to-end autonomous driving takes the raw sensor data as visual input and yields a control command for the vehicle (Figure 2, b) [20, 28]. Particularly, recent breakthroughs in deep learning and computer vision algorithms and the availability of rich sensor devices have been the primary reasons for automotive researchers to leverage the end-to-end learning approach. The advantage of an end-to-end pipeline over its counterpart is that it directly produces driving actions by unifying perception, localization, planning, and control as one combined machine learning (ML) task. Furthermore, computational efficiency is improved via shared backbones in end-to-end learning, and in this way, potential information loss in intermediate layers is also avoided [20, 29].
2.2 Existing issues
AI approaches, which are currently predominated by deep learning algorithms, have brought considerable improvements to many essential components of autonomous driving technology, including advances in perception, object detection, and planning. As the AI-powered driving systems of vehicles advance, the number of AVs deployed to road networks has proliferated significantly in many developed European countries, the US, and Canada over the last decade [30]. However, the aforementioned road accidents involving such cars have caused public skepticism, and many studies have attempted to underscore the current limitations and issues with the design, development, and deployment of autonomous cars on roads. For example, Fleetwood [31] has investigated public health and ethical issues arising with the use of autonomous driving. Their study provides an in-depth analysis of the health issues, especially with the Trolley problem examples [32, 33] (hitting a pedestrian on an icy road or a parked car; driving and hitting five people or changing the direction of the steering wheel and hitting an individual, etc.). Another critical aspect of the work is a concern for the potential rights and liabilities of passengers sitting inside an autonomous car (for example, by using such a car, does a passenger agree to face potential risks; does a passenger have the responsibility and liability to protect other road users if an accident happens?). That study concludes with four directions - clear and cross-disciplinary discussions amongst stakeholders, including a driving system’s action planning choice of an algorithm; enhancing society’s knowledge on the issues and limitations of autonomous driving; confirming society’s knowledge on solutions of the current issues and proper use of AVs; and developing faithful, rational, and monitored standards for public health experts’ attention.
Some studies have directly focused on the concept of ethical crashing (i.e., if crashing is inevitable, how to crash?) and the Trolley problem mentioned above. For instance, the Moral Machine experiment [34], a well-known and hotly debated experiment, investigates a general community’s preferences on applied Trolley problems (inevitable accident scenarios with binary outcomes) and states that “these preferences can contribute to developing global, socially acceptable principles for machine ethics.” However, further discussion on this issue condemns this opinion and draws attention to the lack of safety principles [35], which force deeper consideration of such dilemmas [36]. Sohrabi et al. [37] have explored the outcomes of autonomous driving technology on public health in an urban area. They determined that this technology can affect public health in thirty-two directions: more than half of these pathways (seventeen) are negative, eight are positive, and the remaining ones are unsettled. Another similar study led by Martinho et al. [38] has inspected autonomous vehicle-related ethics in both scientific literature and various industry reports in California. According to this survey, both the scholarly literature and industry documents highlight safety and cybersecurity as main issues and underline concerns for moral decision-making algorithms, human-involved control, ethics, design, privacy, accountability, and sustainability, among other concerns. Burton et al. [39] have identified three open problems in the state-of-the-art development of autonomous systems. The first one is the semantic gap that emerges when a thorough specification of the system is not provided to manufacturers and designers. Another identified issue is the responsibility gap, which arises when an accident happens and the responsibility of either an autonomous system or a human is the cause of this accident remains unresolved. Finally, there is the question of who is responsible for compensating the injured during an accident, which precipitates the third issue: the liability gap. That study also shows that the core of these issues is associated with domain complexity, system complexity, and transferring more decision-making functions from humans to autonomous systems. In a relevant case study, authors relate unpredictable urban environments, driving system complexity, and lack of human override in highly autonomous driving to the enumerated core reasons, respectively. Finally, several studies have attempted to elucidate other related issues such as privacy [40, 41], design and implementation issues [42], legal regulation [43], user concerns [44], IoT challenges [45], integration in smart cities [46], as well as management [47] and security concerns [48]. The key findings outlined in the above studies require an understanding of the causes of these issues and intrinsically give the stakeholders the right to ask “why” questions. So, we immediately observe an immense need for explanations about the performance of self-driving cars. Providing explanations of critical decisions can significantly increase the acceptance of AVs by both the transportation jurisdictions and general community. In the subsequent sections, we discuss established standards and regulations for autonomous driving technology and describe the design principles of AI software architecture for it. We then analyze the need for explanations in AVs, emerging from these multidisciplinary points of view.
2.3 AVs regulations and standards
The issues and growing concerns caused by AI systems create the need to scrutinize the regulation of this technology. As a result, public institutions have initiated the development of regulatory frameworks to monitor the activities of data-driven systems at both a country level and internationally. The focal points of these regulations are mainly to protect the stakeholders’ rights and ensure they have control over their data. For example, the General Data Protection Regulation (GDPR) of the European Union (EU) initiated guidelines to promote the “right of an explanation” principle for users, enacted in 2016 and taking effect in May 2018 [49]. Moreover, the EU has a specially defined strategy on Guidelines of Trustworthy AI that has seven essential requirements, namely 1) human agency, 2) technical robustness and safety, 3) privacy and data governance, 4) transparency, 5) accountability, 6) diversity, non-discrimination, and fairness, and 7) societal and environmental well-being; these principles are all to be applied in AI-based product research and development [50]. Similarly, Mohseni et al. [51] have tabled a broad description of ethical AI with a consequent impact on autonomous systems. In this context, AVs also must comply with these principles and provide intelligible explanations to the engaged stakeholders in support of an investigation of road accidents and other critical conditions involving AVs.
Various organizations have recently proposed guidelines on the regulation of AVs to monitor their compliance with law enforcement. NACTO’s (National Association of City Transportation Officials) statement on automated vehicles proposes nine principles to shape a policy on regulation of future generation AVs [52]. Another well-defined guideline, The Research and Development (RAND) Corporation’s principles,
Figure 3: Multidisciplinary perspectives necessitating XAI for autonomous driving
covers promises and issues of AVs and an association of this technology to law and liability issues. Their principles also provide thorough guidance for public regulators to investigate transportation accidents and make safety recommendations (e.g., [53, 54]) to the regulators and manufacturers of AVs such as the American National Highway Traffic Safety Administration (NHTSA), SAE International, Tesla, and Apple. NHTSA of the US Department of Transportation has a specific federal guideline on automated vehicle policy to improve traffic safety [55] In March 2022, NHTSA announced that automobile manufacturers would no longer have to equip fully autonomous cars with manual control elements, such as steering wheel and braking pedals in the USA [56]. The Government of Canada has also recently released their comprehensive federal guidelines on testing and regulations of automated driving systems [57]. Their documentation provides detailed information and a regulatory road map for the relevant organizations on the engagement with government agencies, pre-trial, testing, and post-test considerations of AVs. In another recently adopted regulation, Germany has published an act on operations of driverless cars, particularly relevant to designated areas of public roads [58]. The UK government has also advanced their interests toward regulated and safe autonomous driving, and hands-free driving was expected to be legally allowed there by the end of 2021 [59]. Other developed countries such as Australia [60], and Japan [61] have also recently launched initiatives for trials of autonomous driving technology.
While the regulations have been set out to ensure legislative norms and user demands are met, some standards provide specifications to achieve a high safety level, quality assurance, efficiency, and environmentally friendly transportation systems. The International Organization for Standardization (ISO) has adopted several standards to define the relevant issues on automated driving. Examples include the ISO 21448 [62], which specifies situational awareness standards to maintain operational safety under the “Safety of the Intended Functionality,” and the ISO 26262 [63] standard defined for the safety of electrical and electronic systems in production passenger vehicles, entitled as “Road vehicles – Functional safety.” In this context, ISO/TC 204 [64] is the primary standard that provides a comprehensive guide on the overall system and infrastructure aspects of intelligent transportation systems (ITS), supporting the standardization of AVs technology. Motivated by ISO/TC 204, some regional initiatives have also imposed relevant standards on the regulation of autonomous vehicles. For instance, The European Committee for Standardization, together with ISO, has the CEN/TC 278 [65] standard that develops acceptable levels of quality, use cases, and best practices for ITS in Europe. It turns out that AVs, or ITS, as a more general field, involve many multidisciplinary foundations to meet the involved stakeholders, insurance, and law enforcement requirements. Thorough documentation on the details of legislation, regulation, and standardization of automated vehicles can be viewed here [66].
3 Explanations in autonomous driving
3.1 The need for explanations in AVs
The need for explanations in autonomous driving arises from existing issues, established regulations and standards covered in previous subsections, and cross-disciplinary views and opinions of society. At the highest level, the necessity of explanations for AVs can be summarized in terms of four perspectives:
  • Psychological perspective: Traffic accidents and safety concerns remain the main cause of the need for XAI in autonomous driving from a psychological point of view [67].
  • Sociotechnical perspective: The design, development, and deployment of AVs should be human-centered, reflecting the target audience’s needs, and taking their prior opinions and expectations into account [68, 69].
  • Philosophical perspective: Explaining AI decisions can provide descriptive information about the causal history of actions performed, particularly in critical situations [70, 71, 72].
  • Legal perspective: It considers all the above-mentioned factors and incorporates them into general regulatory compliance principles for AVs. A notable example is GDPR’s requirements on explanation provision for end users [49].
Overall, we can conclude that the explainability of autonomous driving systems is an expectation and a requirement from a multidisciplinary point of view (Figure 3).
3.2 Potential benefits of explanations in AVs
Considering these multi-dimensional perspectives, explainable autonomous driving can bring the following benefits to the stakeholders:
  • Human-Centered Design
    Getting the end users’ inputs, opinions, and anticipations on the design and development of the semi or fully AVs can help with the acceptance of this technology by the general community [73]. In this context, it is essential to provide user-friendly human-machine interfaces (HMIs) to the stakeholders, such as backup drivers and passengers. Studies show that explanations can be conveyed to end users via light, visual, audio, and textual information using supplied HMIs [74, 75, 76]. For instance, Schneider et al.’s recent empirical study confirms that applying stakeholders’ multi-modal feedback to the simulated design of autonomous driving creates a positive user experience [77]. Thus, a human-centered design meeting users’ needs and expectations is a necessary step for the widespread adoption of AVs technology.
  • Trustworthiness
    As careless and hazardous driving can directly impact the safety of passengers and bystanders, people naturally require an acceptable level of safety of AVs. In addition, understanding the causes of actions or decisions is a natural human requirement. As stated in Riberio et al.’s [78] work, “if the users do not trust a model or a prediction, they will not use it.” Holliday et al. [79] have also empirically shown that providing explanations and perceptible systems significantly increases users’ trust in a system. On the other hand, frequently occurring failures without faithful explanations to stakeholders can seriously damage individual and public trust in intelligent systems. Once trust in an intelligent system is damaged, regaining it can be onerous [80]. In this sense, Israelson and Ahmed [81] have shown that there is an inherent need for algorithmic assurance to build trust in human-autonomous system relationships in their analysis.
  • Traceability
    Explainable intelligent driving systems can help forensic analysts and system auditors understand the entire decision-making process of an autonomous car during the journey via a post-trip analysis. A recorded log data of action-explanation pairs can help understand the car’s temporal actions and shed light on whether particular temporal actions of the car were correct or incorrect. Hence, traceability is an essential requirement of self-driving and can help debug and analyze the existing driving system and provide an opportunity for further enhancements.
  • Transparency and accountability
    If trustworthiness is achieved in intelligent decision-making, it further brings transparency and accountability to AVs technology. Martinho et al. [38] have noted that accountability combines liability and responsibility as a broader concept. In the context of autonomous driving, accountability can be defined as compliance with established legal principles in specific jurisdictions. As the regulatory standards urge the “right to an explanation” as required by GDPR [49], accountability becomes a crucial concept that combines social expectations and legislative norms on the autonomous driving spectrum. In addition, achieving accountability also helps to deal with potential liability and responsibility gaps, as defined by Burton et al., [39], in potential post-accident investigations with the involvement of autonomous cars. Recently Mercedes-Benz has taken a promising step forward and announced that the corporation will take legal responsibility for any accidents that their self-driving systems are engaged in [82]. Mercedes’s declaration of legal culpability is a significant milestone toward the accountability of AVs technology. Hence, explanations promise immense benefits to road users, engaged stakeholders, and AVs manufacturers in general.
3.3 Explanation recipients in AVs
The details, types, and delivery of explanations vary in accordance with users’ identities, technical background knowledge in autonomous driving, and their various functional and cognitive abilities [67, 83]. For instance, a user having little technical expertise on how autonomous cars operate may be satisfied with a simple explanation of a relevant decision/outcome. However, an autonomous systems engineer will need more informative explanations to understand the current functionalities of the car, with the motivation to appropriately “debug” the existing driving system as required. Therefore, the use of domain knowledge and expertise of the explainee is essential to provide pertinent, sufficiently informative, and intelligible explanations [84, 85]. Motivated by a target audience definition of [86] and [67], we can distinguish four groups of the stakeholders in autonomous driving, namely Group 1 - Road users, Group 2 - AVs developers, Group 3 - Regulators and insurers, and Group 4 - Executive management of automobile companies. Figure 4 provides the identity of such stakeholders and their positions in the corresponding classification.
3.4 How to deliver explanations in AVs?
As explainees are classified based on their domain knowledge and needs, explanations and their design and evaluation techniques also vary depending on the context and knowledge of the category of explainees. In fact, explanation construction is one of the major challenges in current XAI research. Zablocki et al. [87] define four “W” questions in XAI-based autonomous driving: 1) Who needs explanations? 2) Why are explanations needed? 3) What kind of explanations can be generated? and 4) When should explanations be delivered? In general, explanations in AI can be distinguished based on their derivation category and classification. Some of the early practical studies have applied explanations to automated collaborative filtering systems [88] and knowledge-intensive case-based reasoning [89]. Another empirical approach attempted to derive explanations based on some intelligibility types [90] and used “why,” “why not,” “what if,” and “how to” type explanations for causality filtering. Furthermore, Liao et al. [91] have interviewed twenty user-interface and design practitioners working in different areas of AI to understand users’ explanatory requirements. By doing so, they have attempted to find the gaps in the interviewers’ products and developed a question bank: the authors represent users’ needs as questions so that users may potentially ask about the outcomes produced by an AI system. Overall, the stakeholder needs-based explanation design can be viewed as one of the promising approaches.
Another popular approach to producing explanations is based on using psychological tools from formal theories, according to the literature review of [92]. Depending on the context and addressee, both explanation derivation methods confirm their usefulness. These explanation generation approaches can find alignment in their application in autonomous driving; since autonomous driving involves people with
Figure 4: Taxonomy of the stakeholders in autonomous driving.
diverse backgrounds in society, relevant XAI design needs inherent adjustments to the context problem. Like their derivation type, explanations also differ depending on the class in which they are included. Through an extensive survey, Omeiza et al. [67] propose the following dimensions of explanations in the context of autonomous driving:
Explanations based on cause filters: Based on available knowledge, explanations use predefined causes to explain the outcome of an event. The explanations are generated based on cause filters such as “why,” “why not,” “how to,” and “what if” queries (e.g., “Why did the car take the left lane instead of the right lane?”). In fact, we can note that this kind of explanations can be used across many autonomous driving operations.
Explanations based on content type: In this category, explanations are classified based on the components or elements involved with the explanations and the way they are presented. Examples of content types include input influence, input sensitivity, case bases, and demographic factors (for instance, explanations based on what input variables (i.e., driving features) contribute more to the predictive actions).
Explanations depending on a model: Here, explanations are distinguished by being either model agnostic or model dependant. For instance, some autonomous driving operations can be condition-specific and some can be general, regardless of the driving conditions (for example, explaining any autonomous driving action, as a model-agnostic rationale, can be specified for this group of explanations).
Explanations based on a system type: This category attempts to capture the properties of the operational system: Omeiza et al. [67] distinguish two kinds of explanations as either data-driven (i.e., explaining the outcome of a predictive model) or goal-driven (explaining an agent’s behaviors based on achieving its goal in a predefined setting).
Explanations with interactivity: Once an explanation is provided, a user may further ask follow-up questions to further understand a provided explanation. This feature brings interactivity into the explanation framework (e.g., a user interface for passengers on board that provides real-time explanations for corresponding actions).
Explanations with concrete scope: This category captures the feasibility and range of explanations that the system can generate by being either local or global. Local explanations are limited to explanations on some or a subset of all possible actions (i.e., explaining a single prediction in a specific traffic scenario). Global explanations, on the other hand, are capable of explaining all high-level decisions from an initial point to the destination, such as why an autonomous car chose a specific map, why it changed the
Figure 5: An example of a CNN for object classification in a real-time traffic scenario.
preplanned route during the travel, and so on.
While Omeiza et al.’s [67] analysis of explanations above primarily focuses on informational content, the effective communication of explanations is also a key factor for good human-machine teaming. In general, the conveyance of explanations to end users is realized through a user interface (UX) or HMI of various types [93]. For instance, an HMI may be an interface to alert the human driver to take over the control of an autonomous vehicle in an emergent situation. Other potential examples are display monitors, sound, light signals, and vibrotactile technology that explain the vehicle’s decision-making intentions and bring situation awareness for people in the loop, as shown in Schneider et al.’s work [77].
4 AI for autonomous driving
Real-time decisions in autonomous driving are based on environmental perception, processing temporal sequence data, and mapping of real-time perception to relevant actions. In this regard, before reviewing relevant XAI techniques, we provide a brief overview of frontier AI approaches - convolutional neural networks and recurrent neural networks with their variations, reinforcement learning, and imitation learning. These AI techniques and their enhancements are dominant techniques that have proven empirical successes in the development of state-of-the-art AVs.
4.1 Convolutional neural networks
Convolutional neural networks (CNN) are an AI architecture to process spatial information, such as images and videos [94]. As a powerful learning technique, CNNs can detect discriminant visual features automatically from an input image and are extensively used for pattern recognition, object classification and detection, and other computer vision applications. CNNs can be regarded as universal nonlinear function approximators. The input x of each layer in a CNN model is organized in three dimensions: width, height, and depth. A typical convolutional neural network is parameterized by a weight vector consisting of a set of weights, W, between neurons and a set of bias values, b:
$\theta=[W,b]$ (1)
During training, a variety of useful and high-level features are extracted in the convolution layer. Then a pooling layer is used to reduce the size of the acquired feature map, to decrease computational costs. After that, the output of these steps is passed to the fully connected layer where neurons, along with the weights and biases, are connected with one another, and a nonlinear activation function is applied to the output of the previous step. Commonly used activation functions with CNN are Sigmoid, Tanh, and ReLu, while ReLu is the most preferred because of its relatively faster convergence. The network further makes
Figure 6: The network structure RNN and LSTM
a final prediction. In the context of autonomous driving, the role of CNN is indispensable for real-time scene understanding tasks, such as object detection, identification, segmentation, and classification. A typical example of CNN architecture for autonomous driving is shown in Figure  5.
While traditional CNNs have been successfully applied to numerous computer vision problems, a limitation of this learning architecture is that it can solely process Euclidean spatial data [95, 96]. However, in real-world applications, there might be a need to capture the spatial relationship between different elements necessitating the processing of non-Euclidean data. In this context, Graph CNNs (GCNNs) [97] have emerged as an alternative to conventional CNNs and are capable of processing any kind of non-grid data by representing elements in a graph structure. In autonomous driving, traffic data such as road networks, traffic flow, and spatial relationships among static and dynamic objects of such a network can be naturally represented as graphs, and GCNNs can effectively capture the dependencies and interactions among these elements. Moreover, as the driving environment is generally dynamic and rapidly changing, an inherent structure of GCNN can allow it to adapt to changes in the environment by dynamically updating the graph structure based on the provided sensor inputs and contextual information. Overall, both traditional CNNs and GCNNs and their augmented variations hold promise for dealing with a variety of challenging vision problems in autonomous driving.
4.2 Recurrent neural networks
Recurrent neural networks (RNNs) are a deep learning architecture for processing temporal and sequential data, such as time series, video, and natural language data [98, 99]. RNNs have a feedback loop to iterate over time phases of sequential data: the output of a previous time step becomes an input to the current step. While iterating over the different time steps, recurrent networks can maintain internal states that contain information about each time step; thus, RNN architectures leverage the concept of “memory,” which uses information from a previous input to yield output in the next time step. A typical RNN architecture has three layers: input, hidden, and output layers. The input layer consists of N units. A sequence of vectors of each time step t denoted as $\{...,{x}_{t-1},{x}_{t},{x}_{t+1},...\}$ is the input of this layer. The input layer is connected to a hidden layer where connections between the units are defined by means of a weight matrix. The hidden units of a hidden layer connect with each other through recurrent connections, and by such a structure, the hidden layer defines the memory of the entire network, formulated as
${h}_{t}=f_{H}({o}_{t}),$ (2)
in which
${o}_{t}={W}_{IH}{x}_{t}+{W}_{HH}{h}_{t-1}+{b}_{h}.$ (3)
The hidden layer is also connected to the output layer with weights ${W}_{HO}$, and based on such a network flow, the units of the output layer are calculated as follows:
${y}_{t}=f_{O}({W}_{HO}{h}_{t}+{b}_{o}).$ (4)
Figure 7: A diagram of reinforcement learning for a self-driving vehicle.
Similar to CNNs, $f_{O}(\cdot)$ is the hidden layer activation function, and ${b}_{h}$ is the bias vector of units of the hidden layer.
A major issue with traditional RNNs is the so-called vanishing gradient problem arising during training: network weights might not be effectively updated if the network is very deep. This may result in considerably small weight values that may reduce the network’s learning ability and force RNNs to have a limited memory capacity. To solve this problem, long short-term memory (LSTM) has been proposed as an enhancement of RNN, which can handle sequential data more effectively and learn better [100]. Compared to a simple RNN architecture, LSTM has “gates” that control the flow of information through the network. Having the ability to preserve long-term memory and retain earlier information in sequence makes LSTMs more practical than traditional RNNs. In general, both RNN and LSTM can be used to predict future position, velocity, and other parameters of autonomous driving. Figure 6 shows a structure of a typical RNN and LSTM.
4.3 Reinforcement learning
Reinforcement learning (RL) is a learning approach where an autonomous software agent interacts with an operational environment and learns to improve its performance via such interactions [101]. RL is a powerful machine learning framework for making real-time and sequential decisions. Generally, sequential decision-making problems are formalized within a setting formally known as Markov decision processes (MDP). An MDP comprises the following parameters:
  • S - a set of states
  • A - a set of actions
  • T- a transition function
  • R - a reward function
  • $\gamma$ a discount factor defined as a fixed value in the (0, 1] interval.
In such a setting, selecting an action ${a\in A}$ results in a new state $s\in S$ with a transition probability ${T(s,a,s^{\prime})\in(0,1)}$, and gives a reward ${R(s,a)}$ to an agent. The goal of an RL agent (i.e., a self-driving vehicle, in our case) is to discover the optimal policy $\pi^{*}$ that results in the maximum expected sum of discounted reward:
[…]
An agent’s reward in starting from a state s taking action a by following a policy $\pi$ is formulated as an action-value function (or Q-function) and defined as follows:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}\Bigg{\{}\sum_{k=0}^{H-1}\gamma^{k}r_{k+1}\mid s_{0}=s,a_{0}=a\Bigg{\}}.$ (6)
Basically, the value function is a measure of how good it is for an agent to be in a particular state [101]. The horizon H is the number of time steps in a given MDP. The MDP setting for RL uses the Markov property: the current state transition depends only on the previous state and previous action. But often, in real-world problems, such as in autonomous driving, an agent might not be able to capture all information and fail to observe all the states of the operational environment. In such a situation, an agent’s interaction with the surroundings is constructed as a partially observable MDP (POMDP). In the POMDP setting, states are replaced by observations [102]. Observations are generated by a latent state, which is not available to an agent in a POMDP. The natural state for a POMDP setting is the distribution on the latent state; this is called a belief state [101]. We can infer that a typical observation is considerably less informative than a natural Markov state, $S_{t}$. So, depending on how perception defines the traffic scenario and tasks, both MDP and POMDP can be used for sequential decision-making in autonomous driving. A general diagram of RL for autonomous driving is described in Figure 7.
4.4 Imitation learning
Imitation learning (IL) is a learning approach mimicking human behavior [103]. Also known as learning from demonstration, the agent learns from a dataset of demonstrations of human experts’ behavior in contrast to an RL agent. The ultimate goal of IL is that the agent imitates the actions of a human expert, and its learned policy $\pi_{il}$ matches the human expert’s policy $\pi_{h}$. There are two well-known types of IL:
1) Behaviour cloning (BC): In this category of IL, a model is trained in a supervised learning manner using a trajectory of state-action pairs $\{(s_{0},a_{0},s_{1},a_{1},\cdots)\}$. The goal is to minimize the loss function that calculates how the IL agent’s actions are different from an expert’s actions. The advantage of behaviour cloning is that is usually simple and efficient and does not require specific reward shaping. On the other hand, misleading correlations between input and output can affect the imitator’s learning ability seriously and lead to compounding errors.
2) Inverse Reinforcement Learning (IRL): Also referred to as Inverse Optimal Control (IOC), IRL learns an underlying reward function that the human expert aims to maximize. Once the agent infers the expert’s reward function, the optimal policy can then be learned via RL. Overall, empirical studies show that IRL can generalize better than BC in unseen states [104]; however, it is computationally costly and more intricate.
5 XAI for autonomous driving: A survey on state of the art
5.1 Previous surveys on XAI for AVs
There exist reviews on XAI for AVs, which provide valuable insights from various perspectives. These studies expose a variety of approaches, from a universal look to an algorithmic point of view. In this sense, the first noteworthy review on XAI for AVs is Omeiza et al.’s work [67]. They study the need for/role of explanations for autonomous driving, and focus on legal requirements, standards, and consumer expectations for the design and development of explainable autonomous driving systems. This provides their basis to present a conceptual XAI framework for modular autonomous driving. In further work, Zablocki et al. [87] present a detailed overview of end-to-end vision-based autonomous driving systems and describe explainability hurdles for AVs from an ML perspective. Finally, in very recent work, Kuznietsov et al. [105] present a systematic review of XAI techniques for modular and end-to-end autonomous driving by focusing on how such techniques improve safety and user trust. In this regard, they propose the SafeX framework for modular autonomous driving by integrating user interface, safety, and explainability.
Table 1: Reviews on XAI for AVs with their focus and coverage
Study Pipelines explored Proposed framework Form and content of explanations? Time granularity of explanations? Emerging paradigms? Methodological taxonomy? Omeiza et al., 2021 [67] modular modular Zablocki et al., 2022 [87] end-to-end Kuznietsov et al., 2024 [105] modular and end-to-end modular Our study modular and end-to-end end-to-end
Our study, as complement, extends the coverage of this previous work in three essential dimensions. First, all three previous surveys specifically focus on form and content of explanations; however, time granularity of explanations has not been investigated. As AVs are real-time decision-making systems, it is crucial to know how explanations must be delivered from the timing perspective. Furthermore, attention-based transformers, large language models, and vision-language models are now at the forefront of AI applied to AVs technologies, and such approaches have not been explored in the aforementioned surveys. Finally, a classification of XAI approaches applied to AVs from an algorithmic/methodological perspective is also a noteworthy nuance missing in these studies. Consequently, our paper extends the above-mentioned reviews by (1) analyzing the temporal sensitivity of explanations, (2) providing a perspective on emerging XAI paradigms, and (3) presenting an algorithmic/methodological taxonomy of XAI methods for AVs. We present the coverage and comparison of these reviews and our work in Table 1.
5.2 Structure of our survey
Our survey presents a comprehensive overview of XAI methods for AVs. In particular, we present a classification of learning approaches in terms of vision, RL and IL, decision trees, logic, user study, and the most recent paradigm shift - large language and vision-language-based explanations for AVs. By adopting insights from the recent industrial trends, emerging AI technologies, and general regulatory-compliance principles, we further present a conceptual framework for explainable end-to-end autonomous driving and describe essential elements of such a framework. Finally, we present a set of promising XAI approaches by describing the missing pieces in the state of the art. We also propose potential solutions with the goal of bridging the gaps and achieving transparency, trustworthiness, and social acceptance in the next-generation AVs.
5.3 Vision-based explanations
As deep neural networks, often in augmented forms of CNNs, power the vision ability of intelligent vehicles, understanding how CNNs capture real-time image segments that lead to the particular behavior of a vehicle is a key concept to achieving visual explanations. In this regard, explainable CNN architectures have resulted in adjustments to generate visual explanations. Zeiler et al. [106] use deconvolution layers to understand the internal representation of CNNs in their seminal work. Hendricks et al. [107] propose a model concentrating on distinguished properties of objects that explain the rationale for the predicted label. Zhou et al.’s [108] saliency map architecture, class activation map (CAM), highlights the discriminative part of an image to predict the label of the image. Moreover, Selvaraju et al. [109] propose an augmented
Figure 8: End-to-end learning of steering angle commands from an input image. Source: [110].
version of CAM, called Grad-CAM, that highlights the derivative of CNN’s prediction with respect to its input. Further examples of backpropagation-based methods include guided-backpropagation, [111], layer-wise relevance propagation [112, 113], and DeepLift [114]. Babiker and Goebel [115, 116] have also shown that heuristics-based Deep Visual Explanations (DVE) provide a justification for predictions of a CNN.
Explaining autonomous driving decisions using visual techniques is also primarily motivated by these studies. Particularly, Bojarski et al.’s work [117] is the first explainable vision approach for self-driving, where the authors propose a visualization method, called VisualBackProp, showing which set of input pixels contributes to a prediction made by CNNs. Their experiments conducted with the Udacity self-driving car dataset on an end-to-end autonomous driving task show that the proposed technique is a useful tool for debugging predictions of CNNs.
Similarly, Hofmarcher et al. [118] propose a semantic segmentation model implemented as a pixel-wise classification that explains underlying real-time perception of the environment. They evaluate the performance of their framework on Cityscapes [119], a benchmark dataset for understanding street scenes. The framework outperforms other popular segmentation models such as ENet and SegNet with 59.8 per-class mean intersection over union (IoU) and 84.3 per-category mean IoU. Interpretability of the model is a plus for unexpected behaviors, allowing to debug the driving system and understand the rationales for temporal decisions of a self-driving vehicle.
Kim and Canny [110] use a causal attention model on top of the saliency filtering that indicates which input regions actually affect the steering control. Their experiments are conducted on the driving datasets - Comma.ai [120], Udacity [121], and Hyundai Center of Excellence in Integrated Vehicle Safety Systems and Control (HCE): This model runs for nearly 16 hours to train CNNs end-to-end from images to steering angles and apply causality filtering to find out which parts of images have high influence in predictions (Figure 8). With this approach, the learned framework provides an interpretable visualization of a vehicle’s actions. As an enhancement of this model, Kim et al. [122] provide textual explanations in their further study. They produce “intelligible explanations” on the decisive actions of a self-driving vehicle using an attention-based video-to-text mechanism and introduce a novel dataset, called Berkeley Deep Drive-X (eXplanation) (BDD-X), that contains annotations for textual explanations and descriptions.
Zeng et al.’s [123] architecture learns to drive an autonomous vehicle safely by following traffic rules, including interaction with road users, yielding, and traffic signals. They use raw LIDAR data and an HD map to generate interpretable representations as 3D detection of objects, anticipated future trajectories, and cost map visualizations. 3D detection instances provide descriptive information so that the model understands the operational environment. Motion forecasting, measured as L1 and L2 distances, explains
Table 2: Explainable vision-based studies for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Bojarski et al.,   [117], 2016 Pixel-based explanations of CNN predictions CNN Visual AVs developers
Kim and Canny  [110], 2017 Explaining behavior of a vehicle controller using heat maps CNN, LSTM Visual AVs developers
Kim et al.,  [122], 2018 Generating textual explanations on a vehicle’s control commands CNN, S2VT, LSTM Visual and Textual All groups
Hofmarcher et al.,  [118], 2019 Visual scene understanding using semantic segmentation Enet, SqueezeNet 1.1, ELU Visual AVs developers
Zeng et al., [123], 2019 End-to-end interpretable neural motion planner FaF, IntentNet Visual AVs developers
Hu et al., [124], 2019 Interpretable multi-modal probabilistic prediction for autonomous driving CVAE, Dynamic time warping, LSTM Visual AVs developers
Xu et al.,  [125], 2020 Explaining object-induced action decisions for autonomous vehicles Faster R-CNN Visual All groups
Kim et al.,  [126], 2020 Advisable learning for self-driving vehicles by internalizing observation-to-action rules Mask R-CNN, LSTM Visual and Textual All groups
Li et al., [127], 2021 Risk object identification via causal inference InceptionResnet-V2, Mask R-CNN, Deep SORT, RoIAlign Textual All groups
Casas et al., [128], 2021 End-to-end model for mapless autonomous driving CoordConv Visual and Textual All groups
Kim et al.,  [129], 2021 Explainable and advisable model for self-driving cars DeepLab v3, Mask R-CNN, LSTM Textual All groups
Wang et al.,  [130], 2021 Enhancing automated driving with human foresight Gaze-based vehicle reference Visual Road users
Chitta et al.,  [131], 2021 Interpretable neural attention fields for end-to-end autonomous driving ResNet, MLP Visual AVs developers
Dong et al.,  [132], 2021 Explainable autonomous driving via an image transformer ResNet-50, Mobilenet-v2, multi-head self-attention Textual All groups
Hanna et al., [133], 2021 Interpretable goal recognition in the presence of occluded factors for autonomous vehicles Goal and Occluded Factor Inference, Monte Carlo Tree Search Visual AVs developers
Mankodiya  et al., [134], 2021 XAI for trust management in autonomous vehicles Random Forest, Decision Tree, AdaBoost Visual AVs developers
Madhav and Tyagi, [135], 2022 Explainable navigational intelligence for trustworthy autonomous driving Grad-CAM, Lime Visual AVs developers
Jing et al., [29], 2022 Interpretable action decision making for autonomous driving Faster R-CNN Visual and Textual All groups
Jacob et al., [136], 2022 Region-targeted counterfactual explanations GANs Visual AVs developers
Zhang et al., [137], 2022 Interrelation modeling for explainable automated driving Faster R-CNN, ResNet-50 Visual AVs developers
Kolekar et al., [138], 2022 Traffic scene understanding via U-Net and Grad-CAM U-Net, GradCam Visual AVs developers
Zemni et al., [139], 2023 Object-aware counterfactual explanations BlobGAN Visual AVs developers
Itkina and Kochenderfer [140], 2023 Trajectory prediction via interpretable self-aware neural networks PostNet Visual AVs developers
Feng et al., [141], 2023 Natural language explanations via semantic scene understanding DeepLabV3 Visual and textual All groups
Hu et al., [142], 2023 Interpretable trajectory prediction and decision-making of AVs LaneGCNN, ResNet Visual AVs developers
Dong et al., [143], 2023 Describing traffic scenes in natural language via attention-based transformer CNN, LSTM, Transformer Visual and textual All groups
Atakishiyev et al., [144], 2023 Explaining autonomous driving actions with visual question answering VGG-19, LSTM, DDPG textual All groups
Feng and Sun [145], 2024 Interpreting self-driving decisions and improving safety by paying more attention to the regions that are near the ego vehicle Multilayer Perceptron, Trajectory-guided Control Prediction Visual AVs developers
Araluce et al., [27], 2024 Using driver attention for an end-to-end explainable decision-making from frontal driving images ARAGAN, MobileNetV2 Visual AVs developers
Echterhoff et al., [146], 2024 Leveraging concept bottlenecks as visual features for predicting control command and explanations of vehicle and human behavior GPT 3.5, Longformer Visual and textual All groups
whether erroneous actions are due to incorrect velocity or calculation of direction. Finally, Cost Map visualization describes the traffic scene via a top-down view. The architecture is evaluated on a large real-driving dataset consisting of 6,500 traffic scenarios with 1.4 million frames and collected across several cities in North America, and measuring traffic rule violation, closeness to human trajectory, and collision. The authors also carry out an ablation study and show the impact of different overrides, input horizons, and training losses on end-to-end learning.
Xu et al. [125] propose object-induced actions with explanations for predictions of an autonomous car. The authors introduce a new dataset called BDD-OIA, as an extension of the BDD100K dataset [147]; this extension is annotated with 21 explanation templates on a set of 4 actions. Their multi-task formulation for predicting actions also improves the accuracy of action selection. The CNN architecture further unifies
Figure 9: Human advice to a vehicle for appropriate action. Source: [126].
reasoning on action-inducing objects and the context of scenes globally. The empirical results of the study on the introduced BDD-OIA dataset show that the explainability of the architecture also enhances action-inducing object recognition, resulting in better self-driving.
In two respective studies, Kim et al., [126, 129] propose an approach that leverages human advice to learn vehicle control (Figure 9). By sensing operational surroundings, the system is able to generate intelligible explanations on the decisive actions (For example, “Slowing down because the road is wet”). The proposed architecture incorporates semantic segmentation with an attention mechanism that enriches knowledge representation. Experiments performed on the BDD-X dataset show that human advice with semantic segmentation and heat maps improves both the safety and explainability of predictive actions of a self-driving vehicle.
As a more recent vision-to-text approach, Atakishiyev et al. [144] employ the visual question answering (VQA) mechanism to explain autonomous driving actions. They train an RL agent and generate driving data showing the self-driving car’s motion from its field of view. They further convert this video to image sequences, manually annotate the images with question-answer (QA) pairs, and encode questions and images with LSTM and pre-trained VGG-19 [148], respectively. The experimental results on five action categories show that VQA is a straightforward, effective, and human-interpretable approach to justify autonomous driving actions. Leveraging frontal images for interpretable decision-making has further been explored by subsequent studies as well [27, 143, 146].
While the mentioned studies focus on vision-based explanations of already obtained predictions of the model, there have been some recent studies paying attention to counterfactual explanations. In the context of automated driving, counterfactual analysis can be described with such an exemplary question: “Given the driving scene, how can it be modified so that the vehicle keeps driving instead of stopping ?” In other words, given the input, counterfactual analysis intends to figure out the distinguished features in this input that cause the model to make a certain prediction by envisioning modification of those features would cause the model to make a different prediction (e.g., Figure 10). Thus, in this case, the predictions obtained by the existing model and the imagined model become contrastive. As the application of counterfactual intervention, Li et al. [127] presents an approach to find out risk objects that result in particular driving behavior. Their method, formalized as a Functional Causal Model (FCM), shows that the random elimination of some objects from the scene changes the driving decision to the contrastive prediction, such as from the “Stop” to “Go” command. In further work, Jacob et al. [136] introduce the STEEX model that uses a pre-trained generative model to produce counterfactual rationales by modifying
Figure 10: An example of a counterfactual explanation generated by STEEX. Graphics credit: [136].
the style of the scene while retaining the structure of the driving scene. Finally, as further enhancement of STEEX, Zemni et al. [149] propose a method called OCTET that generates object-aware counterfactual explanations without depending on the structural layout of the driving scene as backpropagation can optimize the spatial positions of the provided instances.
Overall, we observe a significant focus on perception-based explanations of autonomous driving systems, as such explanations provide an opportunity to better understand how accurately a self-driving vehicle senses the operational environment. We show a summary of vision-based explanations for AVs in Table 2.
5.4 RL and IL-based explanations
Explaining how perceived environmental states are mapped to actions has also recently received attention in the autonomous driving community. In this regard, the field of explainable RL (XRL) is a relatively new and emerging research avenue on XAI [150, 151, 152]. Like vision-based explanations, XRL techniques also aim to provide some forms of justifications on chosen actions of a vehicle either via intrinsically interpretable design or post-hoc explanations. One of the early works in this context is the Semantic Predictive Control (SPC) framework [153], where the authors propose a data-efficient policy learning approach that predicts future semantic segmentation and provides visual explanations of a learned policy. The framework concatenates multi-scale intermediate features from RGB with tiled actions. The joined modules are then fed into the multi-scale prediction model that predicts future features. Finally, in the last part of the pipeline, the information prediction module inputs the latent feature representation and outputs driving signals alongside the semantic segmentation of the scene.
Chen et al. [154] introduce a sequential latent environment model learned with RL and a probabilistic graphical model-based approach that can interpret autonomous cars’ actions via a bird-eye mask. They use video cameras and LIDAR images as input in the CARLA simulator. For the purpose of interpretability of actions and explainability of a learned policy, they generate a bird-eye mask (i.e., Figure 11). Their model outperforms the used baseline models - DQN, DDPG, TD3, and SAC. Similarly, Wang et al. [155] propose an interpretable end-to-end vision-based motion planning (IVMP) to interpret the underlying actions of a self-driving vehicle. They use semantic maps of a bird-view space in order to plan the motion trajectories of an autonomous car. Moreover, the IVMP approach uses an optical flow distillation network that can improve the real-time performance of the network. The experiments conducted on the nuScenes dataset [156] show the superiority of the proposal over modern approaches in semantic map segmentation and imitation of human drivers. In another probabilistic decision-making model, Wang et al. [157] approach lane merging task as a dynamic process and integrate internal states into joint Hidden Markov Model (HMM) and Gaussian Mixture Regression (GMM). The experiments collected on the INTERACTION dataset [158] demonstrate the efficiency of the proposed technique and show that merging at highway on-ramps can be delineated by three interpretable internal states in terms of the absolute speed of a vehicle while merging.
Rjoub et al. [159] have shown that federated deep RL combined with XAI can lead to trusted autonomous
Table 3: Explainable RL and IL-based studies for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Pan et al., [153], 2019 Semantic predictive control for explainable policy learning LSTM, DDPG-SEG, DLA, model-based RL Visual AVs developers
Bansal et al., [160], 2019 Leveraging concept bottlenecks as visual features for predicting control command and explanations of vehicle and human behavior ChauffeurNet, AgentRNN, PerceptionRNN, Imitation learning Visual AVs developers
Cultrera et al., [161], 2020 Explaining autonomous driving by learning end-to-end visual attention CNN, IL Visual AVs developers
Chen et al., [154], 2021 Interpretable end-to-end autonomous driving with latent deep reinforcement learning MaxEnt RL, DQN, DDPG, TD3 and SAC Visual AVs developers
Schmidt et al., [162], 2021 Interpretable and verifiably RL for autonomoous driving learning SafeVIPER, PPO Visual AVs developers
Wang et al., [155], 2021 Learning interpretable end-to-end vision-based motion planning with optical flow distillation IVMP, Optical flow Visual AVs developers
Wang et al., [157], 2021 Uncovering interpretable internal states of merging tasks at highway on-ramps for autonomous driving decision-making GMR, HMM Visual AVs developers
Rjoub et al., [159], 2022 XAI-based federated deep RL for autonomous driving DQN, DQN-XAI, SHAP Visual AVs developers
Renz et al., [163], 2022 Explainable planning for autonomous driving BERT, GRU, Imitation Learning Visual AVs developers
Teng et al., [163], 2022 Interpretable imitation learning for end-to-end autonomous driving Bird’s Eye View model, Imitation Learning Visual AVs developers
Hejase et al., [164], 2022 Interpretable state representation for deep RL in autonomous driving DDQN Visual AVs developers
Cultrera et al., [165], 2023 Visual attention and end-to-end trainable region proposals for explainable autonomous driving IL, Region Proposal Networks, Spatial Transformers Network Visual AVs developers
Shao et al., [166], 2023 Interpretable sensor fusion transformer for safe autonomous driving InterFuser, Imitation Learning, ResNet, GRU Visual AVs developers
Paleja et al., [167], 2023 Interpretable continuous control trees for autonomous driving Differentiable Decision Trees, SAC Visual AVs developers
Kenny et al., [168], 2023 Interpretable Deep RL with Human-Friendly Prototypes for autonomous driving PW-Net, PPO, TD3 Visual AVs developers
Yang et al., [169], 2023 Reward consistency for interpretable feature discovery for autonomous driving PPO Visual AVs developers
Lu et al., [170], 2024 Human-like cognitive maps for enhancing interpretability of autonomous driving Successor Representations, Cognitive Potential Field Visual AVs developers
Liu et al., [171], 2024 Interpretable generative adversarial IL for autonomous driving IL, Signal Temporal Logic Visual AVs developers
Figure 11: RL-based interpretable end-to-end autonomous driving via a bird-eye mask. Credit: [154]
driving. They use a federated learning approach for decision-making and leverage edge computing that enables different devices to train an ML model in a collaborative manner. The model is first developed on the parameter server and further broadcasted to other devices. Then, global ML methods intake updates from these devices regularly, and the process continues until the model performs well enough on driving tasks. Yang et al. [169] have also shown that an intrinsically interpretable RL agent for autonomous driving can be achieved via reward consistency aiming to resolve the gradient disconnection in reward-action mapping.
Finally, within the exploration via interaction context, a few studies have employed various forms of IL techniques for explainable autonomous driving. Cultrera et al. [161] present conditional imitation learning with an end-to-end visual attention model, which identifies those parts of images that have a higher influence on predictions. They test their architecture on the CARLA simulator [172] on four tasks - go straight, turn left, turn right, and follow the lane. Their ablation study focused on box type importance and fixed grid analysis to get an attention map on the images shows that integrated imitation learning and
Table 4: Studies on decision tree-based explanations for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Omeiza et al., [173], 2021 Generating tree-based explanations with and without causal attributions Tree-based representation / User study Textual All groups
Brewitt et al., [174], 2021 Interpretable and verifiable goal recognition with learned decision trees for autonomous driving Decision Tree Visual and Textual AVs developers
Mankondiya et al., [134], 2021 XAI for trust management in autonomous vehicles Random Forest, Decision Tree, AdaBoost Visual AVs developers
Cui et al., [175], 2022 Interpretation framework for autonomous driving Random Forest, SHAP Visual AVs developers
Brewitt et al., [176], 2023 Interpretatable trees for goal recognition in autonomous driving under occlusion OGRIT, Decision Tree Visual AVs developers
attention model enables a car to drive safely and perform relevant maneuvers in real-time.
Leveraging vision for interpretability of an agent’s actions, Teng et al. [177] propose hierarchical interpretable IL (HIIL) technique that unifies bird eye view (BEV) mask with the steering angle to perform actions in complex situations as an end-to-end autonomous driving pipeline. They construct their method as a two-phase task: In the first phase, the pre-trained BEV model is used to interpret the driving environment. Then imitation learning takes the latent features of BEV mask from the first phase and combines them with a steering angle acquired through the Pure-Pursuit algorithm. The experiment performed on the CARLA simulator shows that the proposed method enhances the interpretability and robustness of driving in various circumstances.
Moreover, Renz et al. [163] introduce PlanT, a rigorous IL-based approach that uses transformers for planning. PlanT is able to explain its action decision by recognizing the most important object in its driving segment and outperforms state-of-the-art work on CARLA’s Longest6 Benchmark by 10 points (See [178] for a visual demonstration). In a more recent work in this context, Liu et al. [171] show that combining Signal Temporal Logic with generative IL is also an effective approach for interpretable policy for autonomous driving. Overall, while building interpretable by-design RL or IL agent for autonomous driving is a challenging task, employing external methods, such as logic and vision, can help achieve explainability for the agent’s actions. A summary of RL and IL-based explanations is shown in Table LABEL:tab:rl_il_exp.
5.5 Decision tree-based explanations
Being inherently interpretable and easier to understand a prediction of a model, decision-tree-based explanations have also been investigated in autonomous driving. Decision trees have been proven to describe the rationale semantically for each prediction made by a CNN architecture [179]. Omeiza et al. [173] use decision trees as a tree-based representation that generates scenario-based explanations of different types by mapping observations to actions in accordance with traffic rules. They use human evaluation in a variety of driving scenarios and generate Why, Why Not, What If, and What explanations for driving situations and empirically prove that the approach is effective for the intelligibility and accountability goals of automated vehicles.
Brewitt et al. [174] introduce Goal Recognition with Interpretable Trees (GRIT), a framework that uses decision trees trained from the trajectory data of a self-driving car. The framework, tested on fixed-frame scenarios, is proven empirically verifiable for goal recognition using a satisfiability modulo theories (SMT) solver [180].
Cui et al. [175] use Random Forest for the interpretability purpose on the autonomous car-following task. They employ deep RL for the decision-making of an autonomous car and employ SHAP values to simplify the feature space. Once the agent generates state-action pairs, Random Forest is applied to these pairs and experimental results show the approach works effectively to explain behavior for the designated car-following task. In a recent study, Random Forest has also been proven to detect misbehaving vehicles in Vehicular Adhoc Networks (VANET) in Mankodiya et al.’s work [134]. Thus, being computationally more transparent than traditional deep neural network architectures, decision trees can explain behaviors of a variety of autonomous driving tasks with less computation.
Table 5: Studies on logic-based explanations for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Suchan et al., [181], 2019 An answer set programming-based abductive reasoning for visual sensemaking Answer set programming, YOLOv3, SSD, Faster R-CNN Visual AVs developers
Corso and Kochenderfer [182], 2020 Interpretable safety validation for autonomous driving Signal temporal logic Textual AVs developers
DeCastro et al., [183], 2020 Interpreting policies via signal temporal logic for autonomous driving Signal temporal logic, LSTM, CVAE Visual AVs developers
Kothawade et al, [184], 2021 Explainable autonomous driving using commonsense reasoning ASP, s(CASP) Textual Road users
5.6 Logic-based explanations
While the interpretability of a deployed autonomous driving control model has been the dominant direction for research, there have also been attempts to verify the safety of self-driving vehicles with logical reasoning. In this regard, Corso and Kochenderfer [182] present a technique to identify interpretable failures of autonomous cars. They use signal temporal logic expressions to describe failure cases of an autonomous car in unprotected left turn and pedestrian crossing scenarios. For this purpose, the authors use genetic programming to optimize signal temporal logic expressions that acquire disturbances trajectories causing a vehicle to fail in decisive action. The experimental results show that the proposed approach is effective in interpreting the safety validation of a car.
Suchan et al. [181] have developed an answer set programming-based abductive reasoning framework for online sensemaking for perception and control tasks. In its essence, the framework integrates knowledge representation and computer vision in an online manner to explain the dynamics of traffic scenes, particularly occlusion scenarios. The authors demonstrate their method’s explainability and commonsensical value with empirical study collected on the KITTI MOD [185] dataset and the MOT benchmark [186]. Another experimental study leveraging the concept of answer set programming has been carried out by Kothawade et al. [184]: they introduce AUTO-DISCERN, a system that incorporates common sense reasoning with answer set programming to automate explainable decision-making for self-driving vehicles. They test their rules and show AUTO-DISCERN’s credibility in real-world scenarios, such as lane changing and right turn operations, from the KITTI dataset.
5.7 User study-based explanations
Finally, except for practical applications, some investigations involve users in case studies to understand the effective strategies for explanation generation in autonomous driving tasks. The key idea of a user study is that getting people’s input in designated driving tasks can help improve the adequacy and quality of explanations in autonomous driving. Wiegand et al. [187] perform a user study that identifies a mental model of users for determining an effective practical implementation of an explanation interface. The main research question here is to understand what components need to be visualized in a vehicle so the user can comprehend the decisions of self-driving vehicles. The study discloses that combining an expert mental model with a user mental model as a target mental model enhances the situation awareness of the drivers. Furthermore, Wiegand et al. [188] investigate situations, in which explanations are needed and methods pertinent to these situations. They spot seventeen scenarios where a self-driving vehicle behaves unexpectedly. Twenty-six participants are selected to validate these situations in the CarMaker driving simulator to provide insights into drivers’ need for explanations. As a result of the user study, the authors identify six groups to highlight the primary concerns of drivers with these unexpected behaviors, namely emotion and evaluation, interpretation and reason, the capability of a self-driving car, interaction, driving forecasting, and request times for explanations.
Wang et al. [130] propose an approach that enables a human driver to provide scene forecasting to an intelligent driving system using a purposeful gaze. They develop a graphical user interface to understand the effect of human drivers on the prediction and control of an intelligent vehicle. A simulator is used to
Table 6: Studies on user study-based explanations for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Wiegand et al., [187], 2019 Explaining driving behavior of autonomous cars User study Textual Backup drivers
Wiegand et al., [188], 2020 Understanding situations that a driver needs explanations User study Visual All groups
Wang et al., [130], 2021 Enhancing automated driving with human foresight User study Visual Backup drivers
Schneider et al., [189], 2021 UX for transparency in autonomous driving UEQ-S, AVAM (User study) Visual, Textual, Light All groups
Schneider et al., [77], 2021 Increasing UX through different feedback modalities UEQ-S (User study) Visual, Textual, Audio, Light, Vibration All groups
Shen et al., [190], 2022 Identifying which scenarios need explanations in autonomous driving Friedman test, Pearson correlation, Point-Biserial Correlation Visual Road users
Schneider et al., [191], 2023 The role explanatory information in failure situations in highly autonomous driving UEQ-S, AVAM Visual, Textual All groups
Kim et al., [192], 2023 Timing perspective and mode of explanations for road users in autonomous driving GradCam, Head-mounted display, Windshield display Visual Road users
test and verify three driving situations where a human driver’s input can improve safety of self-driving.
Apart from these works, Schneider et al. involve human participants in their empirical studies to understand the role of explanations for the public acceptance of AVs [77, 189]. They explore the role of explainability-supplied UX in AVs, provide driving-related explanations to end users with different methods, such as textual, visual, and lighting techniques, and conclude that providing context-aware explanations on autonomous driving actions increases users’ trust in this technology. Their subsequent study also confirms that driving explanations can help mitigate the negative impact of AVs failures on users [191]. Finally, Kim et al.’s user study [192] confirms that humans do not need explanations seamlessly, and presenting explanations only in critical driving conditions is preferred to enjoy the trip with an autonomous car and prevent information overload. We summarize user studies on XAI for AVs in Table LABEL:tab:_user_study_exp.
5.8 Large Language Models and Vision-Language Models-based explanations
While the preliminary studies and further works on explainable autonomous driving primarily focused on a combination of various AI techniques revisited above, large language models (LLMs) and vision-language models (VLMs) have recently emerged as a novel paradigm to interpreting AVs decisions and describing traffic scenes. Built on top of Foundation Models, such as GPT [193] and BERT [194], there has been significant progress on building both domain-agnostic and domain-specific LLMs (e.g., GPT-3 [195], GPT-4 [196], LLAMA [197], LLAMA-2 [198], Vicuna [199], Alpaca [200], Claude [201]) and VLMs (e.g., Flamingo [202], LLaVa [203], PaLM-E [204], Video-LLaVA [205], Video-LLaMA [206], Gemini [207], Claude 3 [208]). In this sense, there have been tremendous efforts to build language and vision-language models on top of these base models for interpretable autonomous driving. Overall, based on the recent trend in leveraging these large models for the interpretability purpose in autonomous driving, we observe the following directions:
Presenting live natural language explanations during the trip: The promising work in this context is Wayve’s LINGO-1 [209] architecture. LINGO-1 is a vision-language-action (VLAM) model that provides live natural language explanations for describing a vehicle’s chosen actions in end-to-end autonomous driving. Trained on diverse multimodal (vision and language) datasets, LINGO-1 can describe action decisions and causal factors inducing these actions. The advantage of LINGO-1 architecture is that its explanations are concise, informative, and reflect temporal changes in the driving environment. An example of LINGO-1’s live and temporal explanations is shown in Figure 12. The Wayve team has also achieved live linguistic explanations for autonomous driving in a simulation environment [210].
Video Question Answering as a reasoning technique: An essential characteristic of modern AVs is the
Table 7: Large language and vision-language based explanations for AVs
Study Target audience
Task Algorithms/Methods Delivery format
Dewangan et al., [211], 2023 Language-augmented Bird’s-eye View Maps for Autonomous Driving GPT-4, GRIT Visual, Textual All groups
Xu et al., [212], 2023 VQA and natural language-based explanations for autonomous driving LLAMA 2, CLIP Visual, Textual All groups
Marcu et al., [213], 2023 Video question answering for autonomous driving Vicuna-1.5-7B, GPT-4 Textual All groups
Chen et al., [210], 2023 Improving context understanding in autonomous driving with object-level vector modalities and LLM GPT 3.5, PPO Visual, Textual All groups
Fu et al., [214], 2023 Understanding traffic situations in a closed loop GPT 3.5 Textual All groups
Mao et al., [215], 2023 Interpretable motion planning as language modeling GPT 3.5 Textual Road users
Sha et al., [216], 2023 LLM as a decision-maker in complex driving scenarios ChatGPT, MPC Visual, Textual AVs developers
Wayve Team [209], 2023 Providing live explanations in natural language Integrated vision, language and action architecture Textual All groups
Nie et al., [217], 2023 Interpretable reasoning in complex driving situations in autonomous driving GPT-4, MLP, ViT-G/14 Textual All groups
Park et al., [218], 2024 Video question answering for traffic scene understanding Video-LLAMA, GPT-4 Textual All groups
Wen et al., [219], 2024 LLM-based knowledge-driven approach for interpretable autonomous driving Out-of-box LLM Visual, Textual AVs developers
Yuan et al., [220], 2024 Retrieval-augmented VLM for explainable autonomous driving LanguageBind, MLP, Vicuna-1.5-7B Textual All groups
Chi et al., [221], 2024 GPT-aided explainable decisions for autonomous vehicles GPT, Graph of Thoughts Textual All groups
Atakishiyev et al., [222], 2024 Robustness of a transformer-based VideoQA model against human-adversarial questions and its safety implications for self-driving Video-LLaVA Textual All groups
Figure 12: Live explanations for temporally evolving scenes by LINGO-1. Graphics adapted from the original source: [209]
consideration of human factors in the design and development of this technology and having effective human-machine alignment for trustworthy autonomous driving. In this sense, it is crucial that users-in-the-loop have some form of interaction with AVs during a journey. Motivated by this concept, some recent works have approached conversational user interface between people on board and AVs as a Video Question Answering (VideoQA) task [212, 213, 218]. Asking questions about the behavior of an autonomous system is a part of our intuition, and in the context of AVs, getting answers to traffic-related situations and autonomous car action-related questions can help users have a comfortable and reliable journey. Other practical applications of LLMs and VLMs are interpretable motion planning [215], chain-of-thought-based reasoning for control and decision-making [216, 219], action justification with a control signal [220], predicting the intent of other traffic actors via visual reasoning and cues [211], and understanding the role of video transformer based-explanations on safety of autonomous driving [222] (see Table LABEL:tab:llm_vlm_exp for more details on LLM and VLM-based explanations).
Overall, as an emerging AI technology, LLMs and VLMs have tremendously benefitted AVs from the interpretability aspect, as described in the above studies. However, it is also worthwhile to mention that there are still spaces for improvement of these models as fictitiously generated explanations (e.g., hallucinations) may have serious safety implications and high-stakes consequences for self-driving actions and human life. We describe such caveats and potential solutions in Subsection 7.5 as future work.
Based on a high-level overview of all these studies, driving explanations are generally multi-modal, context-dependent, and task-specific. Moreover, end-to-end learning has become even more popular for highly autonomous decision-making owing to the combination of powerful deep-learning approaches and the availability of advanced sensor devices. As self-driving decisions directly impact road users, the concept of explainability should be accompanied by the safety standards and principles established by transportation regulators. Motivated by this proposition, we can define explainable autonomous driving as follows:
Explainable autonomous driving is a self-driving approach powered by a compendium of AI techniques 1) ensuring an acceptable level of safety for a vehicle’s real-time decisions, 2) providing explanatory information on the action decisions in critical traffic scenarios in a timely manner, and 3) obeying all traffic rules established by the legal entities and regulators.
Driven by this definition and the state-of-the-art works in the above sections, we present a general and conceptual XAI framework for end-to-end autonomous driving and show the necessary components and process steps to achieve regulatory-compliant next-generation AVs.
6 An XAI framework: integrating end-to-end control, regulatory compliance, and explanations
We propose a unifying framework in which methods for developing XAI, end-to-end learning, and safety components are combined to inform general processes of regulatory principles. Each of these components has a concrete role in our framework. In our recent study [15], we have covered a brief description of end-to-end learning for AVs. We extend the scope of that work and describe the essential elements of end-to-end autonomous driving, and the role of and potential challenges with explanations in such a setting. We describe these individual components as follows:
1. An end-to-end control component: Given all possible instances of environment,
$E=\{e_{1},e_{2},...e_{n}\},$
and a compendium of actions
$A=\{a_{1},a_{2},...a_{n}\},$
an autonomous car can take, the overall role of a control system is to map the perceived environment to corresponding actions:
$C:E\mapsto A.$
This mapping intends to ensure that a controller maps the environment to a relevant action of an autonomous system. A control system $C$ is an end-to-end control system (eeC), if $C$ is a total function that maps every instance of an environment
$e\in E$
to a relevant action
$a\in A.$
The two most prevalent learning paradigms for end-to-end autonomous driving are RL and IL [20]. Typically, RL approaches map sensor information to states and produce control signals for an autonomous car. On the other hand, in IL-based end-to-end learning, an agent is trained by imitating an expert’s behavior to learn the optimal policy. In practice, both types of IL - behavior cloning and inverse RL - have been applied to various end-to-end driving tasks as described in the above-mentioned studies.
Figure 13: A diagram of the proposed explainable end-to-end autonomous driving framework.
2. A safety-regulatory compliance component: The role of the safety-regulatory compliance component, srC, is to represent the function of a regulatory agency, one of whose main roles is to verify the safety of any combination of $eeC$ with autonomous vehicle actions $A$:
$srC=f(eeC,A).$
This requirement could be as pragmatic as some inspection of individual vehicle safety (for example, verifying basic safety functions of an individual vehicle for re-licensing). That said, this concept should be considered as a thorough compliance testing of $eeC$ components from vehicle manufacturers to certify their public safety under international and/or national transportation guidelines such as [49] and [57]. The general principles for acceptable functional safety of road vehicles are defined by the ISO 26262 standard [63]. According to this standard, there should be a safety certification development with evidence-based rationales: the vehicle should be able to meet the established functional safety requirement in its operational context. Part 6 of the ISO 26262 standard [223] is dedicated to end-product development for automotive applications within the software level. This guideline includes the design, development, testing, and verification of software systems in automotive applications. Based on these standards, there seem to be two fundamental approaches to confirming regulatory compliance, which we label confirmation of compliance by “simulation,” and confirmation of compliance by “verification,”. These steps are aligned with our observation regarding the role of XAI in confirming regulatory compliance. In the case of the process of establishing regulatory compliance by simulation, the idea is that a selected set of autonomous actions can be simulated, and then assessed to be satisfactory. This approach is perhaps the most familiar, as it arises naturally from an engineering development trajectory, where the accuracy of simulators determines the quality of compliance (e.g., [224]). The confidence of the established compliance is a function of the accuracy and coverage of the simulation. However, this compliance process can be very expensive and prone to safety gaps, especially when consensus on the properties and scope of a simulation is difficult to achieve. Thus, in general, the simulation part can be considered a ”driving school” for AVs: The designed and developed learning software system should be tested rigorously in this phase before such an autonomous system, as a holistic architecture, is deployed to a real vehicle in the physical environment and real roads.
The alternative, verification, is aligned with our own framework and has significant foundational components established in the discipline of proving software correctness, with a long history (e.g., [225]). The general idea is that offline simulation-based autonomous driving is validated on real roads on real AVs via real sensor suites and a learning software stack by passing the safety checks of regulatory compliance.
In addition to safety assurance, another critical requirement of AVs is their ability to defend against adversarial attacks. The ISO/SAE 21434 standard has defined guidelines for cybersecurity risk management for road vehicles, and AVs must also comply with these requirements [226]. As AVs increasingly rely on their automation ability, it is vital that ML software of an intelligent driving system and built-in interfaces can detect and defend against potential cyber-attacks of the broad spectrum, such as electronic control units (ECU) attacks, in-vehicle network attacks, and automotive key-related attacks [227, 228, 229].
We can expect that the potential evolution of the $srC$ processes will ultimately rely on the automation of regulatory compliance testing against all $eeC$ systems. The complexity of $srC$ systems lies within the scope of the testing methods established in a legal framework, where these methods are the basis for confirming a threshold of safety. For instance, a regulatory agency may require at least 90% regulatory-compliant performance of any particular $eeC$ from $N$ safety tests to be performed. However, as a general requirement, this performance must meet ISO 26262 and ISO/SAE 21434 standards to ensure that an autonomous car’s decision-making procedure is aligned with its underlying ML software: The safety features must pass critical checkpoints, and the autonomous car has to have the ability to defend itself against foreseeable adversarial attacks.
3. An explanation component: This constituent of the framework provides understandable insights into the real-time action decisions made by autonomous driving, complying with and corresponding to an $eeC$ and a $srC$. The explanation component must justify how the autonomous car chooses actions along the trip and has to be able to communicate these pieces of information to the relevant users both during the journey and via a post-trip analysis. As analyzed in the reviewed studies, explanations can be described in visual, textual, feature importance format or in hybrid, multi-modal ways and conveyed via light, audio, vibrotactile, and in other forms as needed.
Temporal granularity and conveyance of explanations: While the format and content of explanations have been the primary focus of XAI research, it is noteworthy to underscore that another important consideration, the time granularity of explanations, has not been well-studied in the state of the art. In general, the timing perspective of AVs explanations can be analyzed in three ways: 1) Should explanations be delivered before action is chosen or after action is performed? 2) What is the appropriate lead time for a safe transition from an automatic mode to a human takeover? and 3) Should explanations be delivered seamlessly or only when it is required? We analyze these nuances separately as follows:
1) Timing mechanism of explanations: Delivering timely explanations can help human drivers/passengers react to emergent situations, such as takeover requests, appropriately and prevent a potential danger in the vicinity. According to Koo et al.’s study [74], it is favorable to convey explanations before a driving event is about to happen. This concept has further been validated by Haspiel et al.’s user study, and human judgment shows that explanations should be delivered before action is decided rather than after it is performed [230]. This judgment makes sense as on-time communication of explanations can bring situation awareness for people on board and enable them to monitor an autonomous car’s subsequent action. If the action to be performed soon is hazardous, a human driver or passenger can manually intervene in the situation with such explanations and prevent a potential danger ahead.
2) The impact of lead time on the safe transition from an automated mode to a human takeover: Another important criterion is determining the amount of time needed to alert human actors for a takeover request. In the user study measuring the impact of 4 s vs. 7 s as the lead time on takeover alert, Huang and Pitts [231] show that a shorter lead time leads to a faster transition to human-controlled mode but also lacks the quality of takeover as lack of time may be stressful for a human actor in such situations. A similar conclusion has been acquired by Mok et al. [232] in the case of 2, 5, and 8 s transition times. Wan and Xu [233] have further verified that an insufficient amount of lead time, such as 3 s, results in an impaired takeover performance, and drivers perform better when enough time, such as more than 10 s, is allotted for takeover requests. In general, it can be concluded that lead time for explaining emergent situations to a human and transitioning control should happen within a few seconds, while for non-critical situations, such as post-trip analysis, the amount of time may be as long as needed.
3) All-time or only necessary explanations?: It is also important to consider that humans need to enjoy their trips with AVs and get information from a vehicle when it is necessary. This aspect also applies to the delivery of explanatory information to end users. When the passengers/human drivers are provided with tons of information during the trip, it can lead to mental overload for them [188]. Consequently, it is generally favored that driving decisions and traffic scenes may be described to humans on board when traffic conditions are critical, and people need to be alerted.
It is also noteworthy to specify that AVs must be equipped with need-based HMIs to deliver explanations. There are some challenges with effective automotive HMI design. First, people may have different choices or preferences for HMI (i.e., display monitor, alert interface, etc.). Furthermore, users’ various cognitive and functional abilities must be a crucial factor in the design of user interfaces [83]. For instance, people with visual or hearing impairments may need a customized HMI. Hence, automotive manufacturers must consider the diversity of users, contemplate the timing perspective of HMI explanations in line with relevant actions, and reach a consensus on the best practices with effective HMI design for AVs [234].
Based on the mentioned process steps and crucial elements, we see that achieving the interpretability of self-driving models is challenging, necessitating integration of those steps and cooperation between users and AVs. Consequently, while we argue that transparent and highly autonomous driving is feasible, human factors must be a vital consideration in the design and development of such systems. A simple graphical illustration of our proposed framework with its elements can be seen in Figure 13.
Moreover, through the extensive analysis of the reviewed studies, a holistic view suggests that automotive researchers and practitioners have investigated the safety and explainability aspects of AVs separately in general. Nevertheless, as autonomous driving is a safety-critical application of AI, these concepts must be explored in tandem, and explainability-guided safety improvement deserves more attention. While the majority of AVs’ explanations are post-hoc and provide an opportunity to improve the existing driving system via a retrospective analysis, there is also a need to develop intrinsically interpretable driving systems. One of the seminal studies on the latter topic is Rudin’s [235] work. The author analytically and empirically proves that instead of interpreting black-box ML models in a post-hoc manner, interpretable by-design models should be a preferred choice to design trusted autonomous systems. The author also justifies that the adopted trade-off concept between accuracy and model interpretability is not valid across all domains and domain-specific tasks, such as in computer vision problems. It is possible that an interpretable model can also lead to the same accuracy as the black-box model. From this perspective, we present potential XAI methods with examples and show the role of both live and post-hoc explanations for safe and trustworthy autonomous driving.
7 Mind the gap: Future perspective of XAI in autonomous driving
In this section, we propose potential XAI approaches that can bridge the gap between missing pieces in ongoing research on explainable AVs. We specifically show the current limitations and problems with modern autonomous driving and present guidelines to make AVs more transparent, safer, and robust.
7.1 Trustworthy and Debuggable Explainable Vision
7.1.1 Explainable vision through post-hoc explanations
Explainable vision-directed actions for AVs are based on how high-level features are used to detect and identify objects. The literature reviewed in Subsection 5.3 presents promising attempts to interpret CNN predictions. It is also important that these explanations help debug, find flaws, and improve the existing vision system. We show the importance of vision-based post-hoc explanations in autonomous driving with two examples. As an alternative to the Trolley problems, International Telecommunication Unit (ITU) recently initiated the “‘Molly problem” defined as “A young girl called Molly is crossing the road alone and is hit by an unoccupied self-driving vehicle. There are no eye- witnesses. What should happen next?” [236]. This is where the value of the post-hoc explanation strategy emerges: The vision system of an automated vehicle could provide a rationale on how it perceived, identified, and distinguished on-road objects and why it continued to drive, resulting in hitting the pedestrian. Such explanations are important for transparency in post-accident investigations under regulatory compliance. The second example is the hacking of Tesla’s Autopilot by the McAfee Advanced Threat Research team [237]. The team added a sticker to the label of the actual speed limit, 35 mph, and caused the car’s heads-up display to perceive it as 85 mph (see Figure 14). The car wrongly accelerated beyond the allowed limit in that traffic area.
Figure 14: Extending the middle of 3 by a sticker on the speed limit sign (left side) causes Tesla’s ADAS (right side) to read the limit as 85 mph [237]. The red circle on the right side has manually been added to show the causal effect of the performed hacking.
Even without a careful look at the modified speed limit, humans will not immediately understand why the car sped up in this example. While this is an intentional test performed by humans, similar confusion of the intelligent driving system may be caused by natural phenomena, such as adverse atmospheric conditions. For example, assume that the part of the speed limit sign is covered by muddy rain, and the vehicle’s ADAS perceives 3 as 8, increases speed, and potentially causes a traffic accident. So, we see the importance of a post-hoc explanation once again. “Speed limit shows 85 mph, accelerating” would be a simple, timely description to understand the reason for speeding up in case of a regulatory investigation. So, it turns out that the history of descriptive natural language generated along with each relevant action could be helpful to produce reliable post-hoc explanations for critical traffic scenarios. This technique is helpful to debug and improve the existing system as well.
7.1.2 Explainable vision through intrinsic and live explanations
As explained in Challange 3 in [235], developing interpretable models that concurrently provide explanations is a promising approach towards achieving a transparent vision system, particularly in object classification tasks. Initial specifications on this perspective use a concept of the prototype layer as an addition to a deep neural network: parts of a visual object are decomposed into pieces, and the prototype layer picks out some representative parts of the object during training. When given a new image during testing, the network tries to discover the similarities between those parts of the new image and the ones learned as prototypes in training. In this way, the deep network accumulates evidence from the prototypes and classifies the objects accordingly. Empirical studies of this perspective on classifying handwritten digits, cars, Fashion-MNIST dataset [238], and bird species and cars [239] report nearly the same accuracy as the original black-box models on which they were built. This simple technique inherently produces an explanation in parallel with a prediction of the neural network and does not require any further post-hoc explanation. In the context of autonomous driving, generating live explanations can contribute significantly to accident prevention. Assume that a self-driving vehicle has an in-vehicle person (i.e., a backup driver or a passenger). The vehicle provides a control (i.e., stop) button for emergency use. The built-in interface inside the vehicle shows there is no human ahead crossing the road and continues driving while there is a human ahead actually (i.e., a vision system malfunction). By noticing such an anomaly on time, the person on board can use the emergency button to slow down and/or stop the car and
Figure 15: Model-based RL vs. model-free RL from explainability perspective.
prevent the accident. This simple example shows that live explanations have potential use in autonomous driving and provide an opportunity for safe navigation of a vehicle.
7.2 Explainable Actions Through Model-based Reinforcement Learning
Characterizing AV actions as a sequential decision-making process along the trajectory, such exploration is typically performed within MDP. The most commonly explored approaches for AVs include three types of MDP: IL, model-free RL, and model-based RL, as reviewed in the studies above. IL intends to mimic the behavior of a human driver; this learning process is computationally expensive as it first must gather real-world driving data as training data [240]. Hence, driving policies obtained under this setting can not be controlled at test time [241]. Model-free RL algorithms learn by sensing the environment directly and do not have access to the transition dynamics (i.e., prior knowledge) of the explored environment [101]. Such exploration lacks the explainability of a learned policy: A model-free RL agent explores the driving environment without specific guidance if applied in autonomous driving. It may take a long time to learn an optimal driving policy. This problem is directly addressed in model-based RL: An advantage of model-based RL is that an agent learns the model of the environment first and then adjusts its learning policy according to the dynamics of the environment [242, 243, 244]. This kind of targeted exploration is typically called planning, which inherently makes the learning procedure explainable (see Figure 15). The idea of planning in RL is vital for proper decision-making and has been investigated in detail in the Dyna architectures [245, 246, 247]. The Dyna and its variation, the linear-Dyna architecture, concurrently learn the world model while learning the optimal policy through interacting with the world. Dyna’s planning process creates a predicted future trajectory from an initially provided imaginary state. Based on this structure, model projection generates an optimal action and simultaneously produces a predicted state and a predicted reward. Those last two components can then be visualized and analyzed and help us understand why the agent prefers selecting a particular action at a particular time step as a basis for explanation. As each sequential action of autonomous driving may need a causal explanation, the Dyna architecture and model-based RL, in general, can provide tremendous benefits with their explainability functionality.
7.3 Predictive Knowledge as Knowledge Representation
One of the essential steps for safe driving is to represent sensor-based knowledge accurately. For example, assume an autonomous vehicle approaches a four-way intersection. Based on the interaction with the operational surroundings, the intelligent driving system may take the “full stop,” “turn left,” “turn right,” or “go straight” actions. As prior knowledge, these four predictions can be considered as a set of possible actions that the agent (i.e., the car) may choose. This predictive approach to capturing domain knowledge is regarded as predictive knowledge in RL, and the idea has acquired growing attention in robotics and autonomous systems research [248, 249, 250, 251]. Moreover, such knowledge is also a reliable basis for explanations of the potential actions taken by an autonomous system. In this context, general value functions (GVFs) provide preliminary techniques for representing predictive knowledge. By definition, GVFs aim to capture long-term predictive summaries about actual observations of an RL agent [251, 252]. For instance, an RL agent’s actions for autonomous driving may be judged with questions, and GVFs can be used to represent the corresponding answers. Examples of questions are “What is the probability I will not face a red signal light in the next intersection?” or “What is the expected time till I arrive at my destination, given my current driving policy?” Preliminary experiments performed on robots have demonstrated GVFs’ representation value in learning, and their computational accuracy as a proof of concept has been established [251]. Moreover, Kearney et al. [253] have deeply investigated the role of GVFs from an explainable AI perspective and intuitively analyzed that the question and answer parameters of a GVF can be interpretable. Once an agent inspects its own experience and interprets the decisions to itself, it can also provide an explanatory basis to others through computational estimates. Recent applications of predictive knowledge, formalized as GVFs, on perception problems [254], and learning policies in the real-world autonomous driving [255] setting show the potential benefits of this concept for autonomous driving tasks. Hence, predictive knowledge computationally formalized as GVFs deserves more attention as a knowledge representation approach in autonomous driving.
7.4 Incorporating Commonsense Knowledge into Actions: Temporal Questions and Robust Question-driven Software Hierarchy
Another significant aspect of the intelligibility of an AV’s actions is strongly related to the proper design of its learning software system. This software system, as an end-to-end framework, must provide a rationale for each action taken at $t_{n}$ and explain how this particular action leads to the appropriate subsequent action in the $t_{n+1}$ time step. We can infer that hierarchical software, corresponding to such principles, is an appropriate structure to support explainable decision-making. Such a structure directly reflects our thoughts while driving. For instance, “Will the traffic light change from green to yellow soon?”, “Do the pedestrians ahead intend to cross the road?” or “Why is the car ahead changing its lane?” are just some representative questions that mirror our driving-related considerations while in motion. With this intuition, we can say that the hierarchical software system of an autonomous car can benefit from question-answering-based reasoning, such as using VQA and VideoQA, and users-in-the-loop can judge the temporal actions of AVs by asking meaningful questions to the conversational interface and receiving answers to their queries.
A critical challenge of a QA system is its robustness against tricky or adversarial questions. Traditionally, automotive user interfaces are designed to respond to users’ action-reflecting or conventional questions; however, humans can also ask deliberate questions to observe the interface’s response to the tricky questions. For instance, assume in a left turn scenario, a user enquires, “Why is the car turning to the right?” How should the dialogue system respond to such a question? It is also noteworthy to underscore that people can ask these questions unintentionally (i.e., consider people with visual impairment, color blindness, etc.), and the dialogue system must capture the semantics of such question and present a corrective answer, such as “ The car is actually turning left as traffic light allows it,” as a typical example. In this sense, adversarial VQA has been a recent area of interest to improve the robustness of QA models [256, 257]. However, this concept has not been explored well in autonomous driving and is a promising direction to enhance the robustness and faithfulness of automotive UX.
From the RL perspective, a suitable approach reflecting temporal states is the concept of options [258]. Options are a generalization of actions in which an RL agent has a policy for taking action with the terminal state. The well-established option-critic architecture is motivated by the concept of options [259]. That architecture can learn internal policies and terminal states of options and has proven effective in end-to-end learning of options in the Arcade Learning Environment (ALE). An inherent structure of the option-critic architecture makes it suitable for further development for the learning system of AVs. Driving-related questions are often temporal, and new questions can be generated for the subsequent actions just after a few seconds. The time sensitivity of driving decisions changes dynamically in real
Figure 16: Our approach to AV2.0 vs AV1.0, and potential advantages of AV2.0 over AV1.0 in terms of its AI software stack. The image of the autonomous vehicle has been taken from [260].
time and exposes the vehicle to different levels of risk. Naturally, actions with lower risks are preferred. Nevertheless, in terms of time and computation, we need to explore efficiently to assess the risk levels associated with the corresponding actions. It is possible that focusing only on increasing RL reward may not lead to desired actions in the long term. As an example, Zhang et al. [261] have shown that solely considering reward as a “gold” metric and ignoring levels of risk, as in traditional RL, is not always a perfect decision for an autonomous system, and an RL agent may fail to find the optimal policy with such an exploration. In contrast, incorporating different levels of risks with corresponding actions can help discover an optimal policy in environments through different transition and reward dynamics. Accordingly, a well-constructed question hierarchy and evaluation of risk levels for chosen actions can help achieve timely and informative explanations for AVs in critical traffic circumstances.
7.5 Toward AV 2.0: Unifying vision, language, and action within Embodied AI for safe and explainable end-to-end autonomous driving
While the above subsections describe the potential XAI approaches for vision and action modules individually, we also need to consider AVs’ learning software as a holistic driving system. Three decades of research, starting with ALVINN in 1988 [18] and further succeeding with the DARPA Grand Challenge [262], have achieved significant milestones with traditional AI software. However, recent breakthroughs in Foundation Models in terms of LLMs and VLMs motivate a transition to next-generation AVs. This generation of AVs has been referred to as AV2.0 by industry professionals [263, 264, 265]. The proposal is that the availability of integrated sensor suites, computational resources (i.e., GPU, TPU), and deep learning approaches can help AVs navigate via an end-to-end approach through adaptive learning, scaling, and generalization in complex driving environments. The ability to learn continually through interaction with the environment rather than relying on static datasets has resulted in the emergence of a new direction, labelled as “Embodied AI” [266, 267], and AV2.0 can move forward with such a learning approach. Effectively unifying vision, language, and action within Embodied AI can enable an autonomous car to navigate, interpret, and describe its high-level decisions in real-time. However, safety and explainability components of an end-end self-driving architecture must overcome fundamental challenges in AI described below:
Safety: The established guideline on core problems with AI safety [268] underscores five crucial considerations: avoiding negative side effects, avoiding reward hacking, scalable oversight, safe exploration, and robustness to distributional shift. We analyze the implications of these problems for end-to-end autonomous driving as follows:
  • Avoiding negative side effects: Autonomous driving is primarily associated with the ability of a self-driving car to avoid accidents and maintain a safe distance from stationary and dynamic objects along the planned motion trajectory. However, the scope of the problem is not limited to this feature. Consider a scenario where an autonomous car interacts with another two vehicles, V1 and V2, at a specific moment. While aiming to make safe temporal decisions by itself, the autonomous car must also ensure that it does not implicitly enable V1 and V2 to cause an accident at that road segment as a part of vehicle-to-vehicle (V2V) communication. According to [268], one potential solution to this problem is to leverage cooperative Inverse RL [269], where an autonomous system can cooperate with humans, and a human actor can always shut down the autonomous system in case such a system exhibits undesirable behavior. In the context of autonomous driving, this nuance can be related to an AV’s communication with a human-operated vehicle or other remote operator monitoring an AV’s overall driving safety. One of the prominent methods in this context is Sympathetic Cooperative Driving or SymCoDrive paradigm [270], which trains agents not only to achieve safe driving for themselves but also for human-controlled vehicles by promoting altruistic driving behavior in cooperative autonomous driving. As the deployment of AVs on roads is a gradual process, synergy with human-operated vehicles is a viable approach for socially aware and safe navigation.
  • Avoiding reward hacking: Can we ensure that the end-to-end driving system does not shape its dynamic reward function according to what it sees in less dynamic environments and still apply that reward shaping while transitioning to highly dynamic environments? Particularly, as an embodied AI agent with adaptive learning and generalization ability in unseen environments, reward formulation must account for long horizons ahead and should not adjust its goals for short-term safe driving behavior. This topic has recently been well-investigated by Knox et al. [271]. They propose that flaws in reward shaping for RL-controlled autonomous driving can be identified by eight sanity checks: unsafe reward shaping, potential mismatch between people and reward function’s preferences, undesired risk tolerance via indifference points, learnable loopholes, missing attributes, redundant attributes, and trial-and-error reward design. The study discloses that such sanity checks can capture flaws in reward shaping for autonomous driving that can also exist in reward shaping for other tasks.
  • Scalable oversight: Can humans measure whether AVs perform at a human level or better in general in all driving situations, where in specific moments, evaluating the driving behavior of an end-to-end AV performance may be difficult for humans due to some reasons (i.e., refer to the Molly problem above [236]). While being outside of human override, temporarily, for various reasons, can we trust that AVs will behave safely at that moment? Amodei et al. [268] report that a potential solution to this problem may be semi-supervised RL: an agent can see its reward on a small subset of episodes or times steps. While rewards from all episodes are used to evaluate the agent’s performance, the agent can only use that subset of rewards to optimize its performance under this setting.
  • Safe exploration: Can an AV always make safe decisions when it has a binary choice of actions in a specific time interval? For example, an autonomous car may change its predefined route due to traffic congestion; however, the alternative route may have dangerous potholes or other damaged infrastructure that may lead to risky driving while attempting to save time on the trip.
  • Robustness to distributional shift: A well-known problem with AVs is making the transition from a simulation environment to real roads. For instance, RL-based end-to-end autonomous driving with an impressive performance in a simulation environment may not have the same performance when deployed to a physical autonomous car. Filos et al. [272] have investigated this topic and proposed robust imitative planning, a technique for epistemic uncertainty-aware planning. The key idea is that in case the model has great uncertainty in suggesting a safe course of action, the model can achieve sample-efficient online adaptation by querying the expert driver for feedback. Through several experiments and state-of-the-art results, the authors also release CARNOVEL, a benchmark for evaluating the robustness of driving agents with distribution shifts. Such a benchmark may be a significant part of a robust solution for addressing out-of-distribution scenarios.
These problems reflect a broad spectrum of potential safety issues with end-to-end AVs. However, we argue that the proposal misses yet another essential concept, namely fail-safe ability. This concept has been investigated in some recent work [273, 274, 275]; however, the recent proposals of the next-generation AVs do not explicitly consider this functionality as an integral component of this technology. Human drivers often have a rest once they feel tired on long trips, and a short rest may help them feel mentally/physically better in the next phase of their driving. The same example can applied to AVs as well. Due to internal reasons (e.g., temporary system malfunction) or external factors (e.g., extremely adverse weather conditions), AVs may need to pause their trip temporarily and prevent further high-stakes consequences ahead. Such capability should not be considered a limitation of AVs; on the contrary, it is an optimal design strategy that foresees potential issues due to any factors and makes AVs behave safely by directing them to “have a short rest.”
Explainability: The reviewed studies in Section 5 show a significant milestone in the explainability of self-driving systems. However, there are still significant gaps and challenges to achieving accurate and timely explanations in all phases of trips. For instance, as of September 2023, it is reported that LINGO-1 exhibits roughly 60% performance in its natural language and VQA-based explanations compared to human-level performance [209].
Apart from informational content, another critical aspect deserving attention is the timing perspective of such explanations: the lead time for emergent scenarios, perhaps using extensive scenario-based evaluations or case-based reasoning, must be engineered appropriately. Furthermore, a well-known problem with large pre-trained models, hallucinations, is another challenge in explanation delivery. Particularly, in QA models, the model must generate a response based on the joint question and scene-based semantics rather than being influenced by the question itself, such as in the case of adversarial QAs described in the previous subsection. We have recently performed an empirical study [222] on the latter and shown that even advanced VLMs can fail to detect the language bias in QA models and present incorrect explanations in case of human adversarial questions. This issue, in turn, may damage user trust and can also have negative safety implications for self-driving. So, we argue that large pre-trained models’ construction mechanism can be adjusted and regulated with common sense and human-defined concepts, as also posited by Kenny and Shah [276]. Hence, designing robust QA models deserves more attention to enable meaningful and trustworthy dialogues between users and AVs. These features are key for achieving effective human-AI alignment [277, 278], trust [279, 280, 281], and public acceptance [67, 281, 282] with AV2.0. We describe our approach to AV2.0 and its difference from AV1.0 in Figure 16.
8 Conclusion
We have presented a systematic overview of state-of-the-art investigations, emerging paradigms, and a future perspective of XAI approaches for autonomous driving. Insights from these studies reveal the existing gaps, and we have proposed a conceptual framework for explainable autonomous driving by incorporating missing pieces. The key idea is that AVs need to achieve regulatory-compliant operational safety and explainability in real-time decisions with their increasing automation ability. Together with a detailed overview of the state of the art in XAI-based autonomous driving, our work contributes as a cause-effect-solution perspective. We elaborate on the notion of cause by identifying current gaps, concerns, and a variety of issues specified while denoting effect through the public reluctance on the use of autonomous driving at a broader level. We provide a solution through the proposed framework and a set of promising XAI approaches for a future direction. This paper can benefit automotive researchers and practitioners in understanding the emerging paradigms and industrial trends in XAI approaches for autonomous driving and help achieve ethical, responsible, and trustworthy AVs in the next generation.
Acknowledgment
We acknowledge support from the Alberta Machine Intelligence Institute (Amii), from the Computing Science Department of the University of Alberta, and the Natural Sciences and Engineering Research Council of Canada (NSERC). Shahin Atakishiyev also acknowledges support from the Ministry of Science and Education of the Republic of Azerbaijan.